The length of a rectangle is 2 more than half the width. If the perimeter is 2, find the length and width
Answer by checkley77(12844)

L=W/2+2
2L+2W=2
2(W/2+2)+2W=2
W+4+2W=2
3W=2-4
3W=-2
W=-2/3 ANS. FOR THE WIDTH.
********************************
Something is wrong here.
Can't have a negative width.
Is there something missing from the problem?